a father is 5 times as old as his son, his son is 6 years old, after how many years will the father be 4 times as old as his son?
Answer by richwmiller(17219)

5s=f,
s=6,
5*6=30=f
f+x=4(s+x)
30+x=4*(6+x)
30+x=24+4x
6=3x
2=x
check
30+2=4*(6+2)
32=4*8
32=32
ok